OpenStudy (anonymous):

A chef is using a mixture of two brands of Italian dressing. The first brand contains 9% vinegar, and the second brand contains 14% vinegar. The chef wants to make 240 milliliters of a dressing that is 12% vinegar. How much of each brand should she use?

OpenStudy (anonymous):

@dpaInc thank you

OpenStudy (anonymous):

there is going to be two different ammounts one for the first brand and another for the second

OpenStudy (kropot72):

Amount of vinegar in final dressing = 240 * 0.12 mL 0.09A + 0.14B = 28.8..................(1) A + B = 240.................................(2) Where A and B are the amounts of the two brands. Using (1) and (2) solve for A and B

OpenStudy (mertsj):

Let x = amount of 9% dressing to use. Then 240-x amount of 14% dressing to use. .09x+.14(240-x)=.12(240)

OpenStudy (anonymous):

so how exactly do i find the second one again?

OpenStudy (anonymous):

can i set 1 and 2 equal to eachother

OpenStudy (kropot72):

No. What you do is to find an Equation for A and sustitute. A = 240 - B 0.09(240 - B) + 0.14B = 28.8 21.6 - 0.09B + 0.14B = 28.8 0.05B = 7.2 Can you find B now?

OpenStudy (mertsj):

9x+14(240-x)=12(240) 9x+3360-14x=2880 -5x=-480 x=96=amount of 9% dressing to use 240-x=144= amount of 14% dressing to use

OpenStudy (kropot72):

*substitute

OpenStudy (kropot72):

Basically the method used by @Mertsj is an accelerated version of my method. Both methods give the correct answers when correctly used.

OpenStudy (anonymous):

ok so i got for the first one 28.8 and the second one i got 144

OpenStudy (anonymous):

did i use the methods correctly?

OpenStudy (kropot72):

No. The correct answers are 144 mL and 96 mL

OpenStudy (anonymous):

dangit

OpenStudy (anonymous):

144 belongs to the second dressin right

OpenStudy (kropot72):

You correctly got 144 mL for the amount of one of the brands of dressing. Allyou need to do is to subtract 144 from the final amount of the mixture of dressings to find the amount of the other brand of dressing. 240 - 144 = ? mL

OpenStudy (anonymous):

96

OpenStudy (kropot72):

You got it :) 96 mL

OpenStudy (anonymous):

so once again i ask 144 belongs to the first dressing 96

OpenStudy (anonymous):

belongs to the second

OpenStudy (kropot72):

96 mL of the 9% vinegar dressing and 144 mL of the 14% vinegar dressing are the .correct allocations

OpenStudy (anonymous):

once again thanks for the help